Lakeview is 77 miles from Springfield. Irania must go to the law library in Springfield to get a document for the law firm for which she works. Find how long it takes her to drive round-trip if she averages 44 mph.
The round-trip from Lakeview to Springfield takes approximately __ hours and __ minutes. (Round to the nearest minute as needed.)

Answer by checkley79(3341)

2*77/44=154/44=3.5 hours (3 hours & 30 minutes.